# Predicting small molecules solubilities on endpoint devices using deep ensemble neural networks

### v1

@@ 4.3 Adversarial training
Comparing kde4-NoAdv and kde4, using adversarial training seems to decrease model performance. It can be seen in Table 1 that using adversarial perturbation increased the RMSE from to and to in solubility challenges dataset 1 and 2_1, respectively. However, the RMSE decreased from to in dataset 2_2. Using adversarial perturbation affected our kde4’s performance by a change in RMSE of .
The inconsistent performance improvement observed by using adversarial training was further investigated with models in which the dataset was augmented. Due to the lack of multiple string representations in the training dataset, it is known that kde4 may have some problems generalizing the learning. A generalization issue could direct the adversarial perturbation in a non-physical direction because the model does not have complete knowledge about the chemical representation space. This hypothesis is reinforced when we compare kde10-NoAdv and kde10. When using adversarial training on a model trained with an augmented dataset, the performance improvement is more noticeable () and consistent for all the test datasets.
@@ 4.4 Deep ensemble size
To investigate the effects of increasing the ensemble size, we trained models with an ensemble of 4, 8, and 10 models. Given the previous results, these models used LSTM as the bi-RNN layer and were trained on the augmented dataset. Specifically for the solubility challenge 2_2, the most complex set to predict, these models presented an RMSE of , , and , respectively. The same order of performance was observed in all test sets (See Table 1), showing that increasing the ensemble size consistently improved performance.
@@
The increase in ensemble size led to a decrease in both uncertainties. AU distributions for the kde4 are centered around 4 logS , displaying a long tail that extends to values as high as 20 logS in the worst case (solubility challenge 2_2). A similar trend is seen in EU distributions. On the other hand, the kde10 model results in narrower distributions. The mean of these distributions remains relatively unchanged, but a noticeable reduction in the extent of their tails can be observed. AU distribution ends in values around 10 logS.
@@ 5 Discussion
| test dataset | SolChal1 | SolChal2_1 | SolChal2_2 | ESOL | ||||||
| Model | RMSE | MAE | RMSE | MAE | RMSE | MAE | RMSE | MAE | RMSE | MAE |
| RF | 1.169 | 0.848 | 1.121 | 0.914 | 0.950 | 0.727 | 1.205 | 1.002 | ||
| DNN | 1.596 | 1.118 | 1.540 | 1.214 | 1.315 | 1.035 | 1.879 | 1.381 | ||
| DNN | 1.641 | 1.187 | 1.261 | 1.007 | 1.371 | 1.085 | 2.189 | 1.710 | ||
| kde4 | 1.788 | 1.389 | 1.273 | 0.984 | 1.137 | 0.932 | 1.511 | 1.128 | 1.397 | 1.131 |
| kde8 | 1.704 | 1.332 | 1.247 | 0.984 | 1.044 | 0.846 | 1.418 | 1.118 | 1.676 | 1.339 |
| kde10 | 1.763 | 1.366 | 1.095 | 0.843 | 0.983 | 0.793 | 1.263 | 1.051 | 1.316 | 1.089 |
| Linear regression[39] | 0.83 | 0.75 | ||||||||
| UG-RNN[41] | 0.72 | 0.51 | 0.90 | 0.74 | ||||||
| RF[37] | 0.93 | |||||||||
| Consensus[55] | 1.140 | 0.778 | 0.91 | |||||||
| SolTranNet[44] | 1.711 | 1.004 | 1.295 | 2.99 | ||||||
| GNN[50] | 1.0722 | 0.7256 | 0.91 | 1.17 | ||||||
| SMILES-BERT[91] | 0.47 | |||||||||
| MolBERT[48] | 0.552 | 0.531 | ||||||||
| RT[45] | 0.0547 | 0.73 | ||||||||
| MolFormer[49] | 0.278 | |||||||||
| SolvBert[92] | 0.47 | 0.925 | ||||||||
@@
The Regression Transformers (RT) validation set includes several regression tasks, not only LogS prediction.
After extensively investigating the hyperparameter selection, we compared our model with available state-of-art models from the literature. Performance metrics on withheld validation data, the solubility challenge datasets, can be found in Table 2. Parity plots for our chosen models are presented in Figure 3.
@@
Focusing on the solubility challenge 1 dataset[25], kde10 is only RMSE units worse than the best model available in the literature[41]. The RMSE of the participants of the challenge was not reported.[27] The primary metric used to evaluate models was the percentage of predictions within an error of 0.5 LogS units (called ). Computing the same metric, kde10 has a percentage of correct prediction of 44.4%, a result better than 65% of the participants. The participant with the best performance presented a of 60.7%.
@@
The architecture of the models was not published in the findings of the first challenge.[27] Nevertheless, the findings for the second challenge[28] investigated the participants more thoroughly. Participants were asked to identify their models’ methods and descriptors used. The challenge is divided into two datasets. Set-1 contains LogS values with an average interlaboratory reproducibility of 0.17 LogS. Our kde10 achieve a RMSE of 0.983 and a of 40.0% in this dataset. Therefore, our results perform better than 62% of the published RMSE values and 50% of the . In addition, the model with the best performance is an artificial neural network (ANN) that correctly predicted 61% of the molecule’s LogS using a combination of molecule descriptors and fingerprints. The second dataset (set-2) contains molecules whose solubility measurements are more challenging, reporting an average error in reproducibility of 0.62 LogS. The kde10 achieves an RMSE of 1.263 and a of 23.3%. It performs better than 82% of the candidates when considering the RMSE. Surprisingly, does not follow this outstanding performance, being more significant than only 32% of the participants. Regarding the literature, kde10 has an RMSE only higher than a GNN that used an extensive set of numeric and one-hot descriptors in their feature vector.[50] Our model performs better than a transformer model that uses SMILES-string and an adjacency matrix and inputs.[44] The performance of those models is available in Table 2.
@@
Notably, all participants in the solubility challenge 2 submitted a kind of QSPR or descriptor-based ML. Using descriptors provides an easy way to ensure model invariance concerning molecule representation and is more informative since they can be physical quantities. However, selecting appropriate descriptors is a crucial step for developing descriptor-based ML models. It often requires specialists with a strong intuition about the relevant physical and chemical properties for predicting the target quantity. Our approach, on the other hand, is based on extracting information from simple strings-representations, a simpler raw data. Furthermore, we could achieve state-of-art performance while balancing the model size and complexity and using a raw input (a simple string) to simplify its usage.
@@
Lastly, transformers models have been used to address the issue of accurately predicting the solubility of small compounds. The typical workflow for transformers involves pre-training the model using a large dataset and subsequently fine-tuning it for a specific downstream task using a smaller dataset. Most existing models were either pre-trained and fine-tuned on the ESOL[39] dataset or pre-trained on a larger dataset and fine-tuned using ESOL. Hence, the generalizability of those models cannot be verified. In a study by Francoeur and Koes [44], they considered two versions of their model, SolTranNet. The first version of SolTranNet was trained with the ESOL dataset using random splits. This approach achieved an RMSE of 0.289. Subsequently, the deployed version of SolTranNet was trained with the AqSolDB[1]. When ESOL was used to evaluate their deployed version, the model presented an RMSE of 2.99. While our model achieved an RMSE of 1.316 on ESOL, outperforming SolTranNet deployed version, it cannot be compared with other models trained on ESOL.
@@ 6 Conclusions
Our model was able to predict LogS values directly from SMILES or SELFIES string representations. Hence, there is no need for descriptors selection and construction. Using only raw data, our model could match state-of-art performance in datasets that are challenging to predict accurately.
In addition, carefully compromising between performance and complexity, we implemented a web application using TensorFlow JS. This application can satisfactorily run on any device with limited computational resources, such as laptops and smartphones. This excludes the need to rely on a server to run the application, improving usability and flexibility and decreasing implementation costs.
7 Data and code availability
All code needed to reproduce those results are publicly available on the following GitHub repository: https://github.com/ur-whitelab/mol.dev. The model is also publicly accessible at the following address: https://mol.mol.dev/.

### v4

@@ 3.3 Adversarial training
Comparing kde4-NoAdv and kde4, using adversarial training decreases model performance. It can be seen in Table 1 that using adversarial perturbation increased the RMSE from to and to in solubility challenges dataset 1 and 2_1, respectively. However, the RMSE decreased from to in dataset 2_2. Using adversarial perturbation affected our kde4’s performance by a change in RMSE of .
The inconsistent performance improvement observed when using adversarial training was further investigated with models in which the dataset was augmented. Due to the lack of multiple string representations in the training dataset, it is known that kde4 may have generalization problems. A generalization issue could direct the adversarial perturbation in a non-physical direction because the model does not have complete knowledge about the chemical representation space. This hypothesis is reinforced when we compare kde10-NoAdv and kde10. When using adversarial training on a model trained with an augmented dataset, the performance improvement is more evident () and consistent for all the test datasets.
@@ 3.4 Deep ensemble size
To investigate the effects of increasing the ensemble size, we trained models with an ensemble of 4, 8, and 10 models. Given the previous results, these models used LSTM as the bi-RNN layer and were trained on the augmented dataset. Specifically for the solubility challenge 2_2, the most complex set to predict, these models presented an RMSE of , , and , respectively. Therefore, increasing the ensemble size consistently improved performance. We also observed this improvement on the other datasets (See Table 1).
@@
The increase in ensemble size led to a decrease in both uncertainties. AU distributions for the kde4 are centered around 4 logS , displaying a long tail that extends to values as high as 20 logS in the worst case (solubility challenge 2_2). A similar trend is observed in EU distributions. On the other hand, the kde10 model results in narrower distributions. The mean of these distributions remains relatively unchanged, but a noticeable reduction in the extent of their tails can be observed. AU distribution ends in values around 10 logS.
@@ 4 Discussion
| SolChal1 | SolChal2_1 | SolChal2_2 | ESOL | |||||
| Model | RMSE | MAE | RMSE | MAE | RMSE | MAE | RMSE | MAE |
| RF | 1.121 | 0.914 | 0.950 | 0.727 | 1.205 | 1.002 | ||
| DNN | 1.540 | 1.214 | 1.315 | 1.035 | 1.879 | 1.381 | ||
| DNN | 1.261 | 1.007 | 1.371 | 1.085 | 2.189 | 1.710 | ||
| kde4 | 1.273 | 0.984 | 1.137 | 0.932 | 1.511 | 1.128 | 1.397 | 1.131 |
| kde8 | 1.247 | 0.984 | 1.044 | 0.846 | 1.418 | 1.118 | 1.676 | 1.339 |
| kde10 | 1.095 | 0.843 | 0.983 | 0.793 | 1.263 | 1.051 | 1.316 | 1.089 |
| Linear regression[25] | 0.75 | |||||||
| UG-RNN[34] | 0.90 | 0.74 | ||||||
| RF w/ CDF descriptors[27] | 0.93 | |||||||
| RF w/ Morgan fingerprints[36] | 0.64 | |||||||
| Consensus[88] | 0.91 | |||||||
| GNN[89] | 0.91 | 1.17 | ||||||
| SolvBert[90] | 0.925 | |||||||
| SolTranNet[41] | 1.004 | 1.295 | 2.99 | |||||
| SMILES-BERT[91] | 0.47 | |||||||
| MolBERT[40] | 0.531 | |||||||
| RT[42] | 0.73 | |||||||
| MolFormer[43] | 0.278 | |||||||
@@
After extensively investigating the hyperparameter selection, we compared our model with available state-of-art models from the literature. Performance metrics on the solubility challenge datasets can be found in Table 2. Parity plots for our chosen models are presented in Figure 3.
@@
Comparing the performance of different models is a complex task, as performance metrics cannot be directly compared across models evaluated on distinct datasets. To address this issue, Panapitiya et al. [89] curated a large and diverse dataset to train models with various architectures and molecular representations. They also compared the performance of these models on datasets from the literature[92, 24, 93, 94, 25, 95, 29, 88, 30, 96, 97, 98]. Although their models achieved an RMSE of on their test set, using descriptors as molecular representations resulted in RMSE values ranging from to when applied to other datasets from the literature. According to their study, the Solubility Challenge datasets by Llinàs et al.[29, 30] were found to be particularly challenging due to their more significant reproducibility error. Therefore, we focused on the Llinàs datasets to compare our performance with the literature.
@@
Focusing on the solubility challenge 1 dataset[29], kde10 is only RMSE units worse than the best model available in the literature[34]. The RMSE of the participants of the challenge was not reported.[31] The primary metric used to evaluate models was the percentage of predictions within an error of 0.5 LogS units (called ). Computing the same metric, kde10 has a percentage of correct prediction of 44.4%. This result would place our model among the 35% best participants. The participant with the best performance presented a of 60.7%.
@@
The architecture of the models was not published in the findings of the first challenge.[31] Nevertheless, the findings for the second challenge[32] investigated the participants more thoroughly. Participants were asked to identify their models’ architecture and descriptors used. The challenge is divided into two datasets. Set-1 contains LogS values with an average interlaboratory reproducibility of 0.17 LogS. Our kde10 achieve an RMSE of 0.983 and a of 40.0% in this dataset. Therefore, our model performs better than 62% of the published RMSE values and 50% of the . In addition, the model with the best performance is an artificial neural network (ANN) that correctly predicted 61% () of the molecule’s LogS using a combination of molecule descriptors and fingerprints. The second dataset (set-2) contains molecules whose solubility measurements are more challenging, reporting an average error in reproducibility of 0.62 LogS. The kde10 achieves an RMSE of 1.263 and a of 23.3%. It performs better than 82% of the candidates when considering the RMSE. Surprisingly, does not follow this outstanding performance, which is more significant than only 32% Regarding the literature, kde10 has an RMSE only higher than a GNN that used an extensive set of numeric and one-hot descriptors in their feature vector.[89] Our model performs better than a transformer model that uses SMILES-string and an adjacency matrix and inputs.[41] The performance of those models is available in Table 2.
@@
Notably, all participants in the solubility challenge 2 submitted a kind of QSPR or descriptor-based ML model. Using descriptors provides an easy way to ensure model invariance concerning molecule representation and is more informative since they can be physical quantities. However, selecting appropriate descriptors is crucial for developing descriptor-based ML models. It often requires specialists with a strong intuition about the relevant physical and chemical properties for predicting the target quantity. Feature-based models are still being considered to be the SOTA of solubility prediction. Recently, studies investigating different descriptors and fingerprints were performed.[36, 99] These studies showed that similarly to the impacts of data quality[28], molecular representation also has a great impact on models’ performance. Despite Tayyebi et al. [36] being able to achieve an MAE of 0.64 on solubility challenge 1 when using Morgan fingerprints (MF), Zagidullin et al. [99] reported poor performance when using MF. Our approach, on the other hand, is based on extracting information from simple string representations, a more straightforward raw data. Furthermore, we could achieve SOTA performance while balancing the model size and complexity and using a raw input (a simple string). This simplified usage enables running the model on devices with limited computing power.
@@
Lastly, transformer models have been used to address the issue of accurately predicting the solubility of small compounds. The typical workflow for transformers involves pre-training the model using a large dataset and subsequently fine-tuning it for a specific downstream task using a smaller dataset. Most existing models were either pre-trained on the ESOL[25] dataset or pre-trained on a larger dataset and fine-tuned using ESOL. Hence, the generalizability of those models cannot be verified. In a study by Francoeur and Koes [41], they considered two versions of their model, SolTranNet. The first version of SolTranNet was trained with the ESOL dataset using random splits. This approach achieved an RMSE of 0.278. Subsequently, the deployed version of SolTranNet was trained with the AqSolDB[1]. When ESOL was used to evaluate their deployed version, the model presented an RMSE of 2.99. While our model achieved an RMSE of 1.316 on ESOL, outperforming the SolTranNet deployed version, it cannot be compared with other models trained on ESOL.
@@ 5 Conclusions
We used TensorFlow.js, the JavaScript library of TensorFlow, to develop a deep ensemble recurrent neural network (RNN) capable of accurately predicting LogS values from SMILES or SELFIES string representations. This model is accessible via a static website at https://mol.dev/. The significant contributions of this research include: (1) Demonstrating the feasibility of using string representations for solubility prediction; (2) establishing that using string representations does not significantly compromise performance, with our models achieving results comparable to state-of-the-art (SOTA) models on the datasets by Llinas et al.; (3) enhancing prediction reliability and practical applicability by incorporating uncertainty estimates; and (4) significantly facilitating model usability through the deployment on a static website, eliminating the need for domain-specific data or expertise.
Our based on a deep ensemble of recurrent neural networks (RNNs) model was trained using SMILES randomization for data augmentation on the AqSolDB dataset and validated using the solubility challenges by Llínas et al. It directly processes molecular string representations, such as SMILES or SELFIES, to predict solubility without relying on pre-selected descriptors. This approach not only simplifies the prediction process but also enhances its applicability across a broader chemical space. In addition, we show that this deep ensemble RNN model could achieve similar performance compared to a random forest (RF) using PaDEL descriptors. RFs with descriptors were shown to perform relatively well in other datasets.
By carefully compromising between performance and complexity, we developed a model with acceptable performance and that is not computationally intensive. It enables us to host the model on a static website using TensorFlow JS. Our model was designed to operate on devices with limited computational resources, aiming to broaden the accessibility of advanced solubility prediction tools. This application can satisfactorily run on any device with limited computational resources, such as laptops and smartphones. This approach ensures wider applicability, catering to the needs of users without access to high-performance computing facilities, improving usability and flexibility, and decreasing implementation costs. We believe this is a considerable step in improving the usability of deep learning models and promoting such models to a broader scientific community.
Data and code availability
All code needed to reproduce those results is publicly available on the following GitHub repository: https://github.com/ur-whitelab/mol.dev. The model is also publicly accessible at the following address: https://mol.dev/.
Authors contribution
M.C.R. implemented the deep learning model, performed the training and hyperparameters optimization, tested the model’s performance, analyzed the results, and wrote this manuscript. A.D.W. idealized the project, proposed the model to be used, implemented the deep learning approach, and developed http://mol.dev/.
Conflicts of interest
The authors have no conflicts to declare.
Acknowledgments
The authors acknowledge the National Institute of General Medical Sciences of the National Institutes of Health (NIH) under award number R35GM137966. This research used the computational resources and structure provided by the Center for Integrated Research Computing (CIRC) at the University of Rochester.